# Quantum Mixing and Benjamini–Schramm Convergence of Hyperbolic Surfaces

### v1

@@ 5. Propagators
We want to show that the following expression goes to 0 as $\varepsilon$ goes to zero:
$\int_{\mathbb{H}}\left|\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)-\chi_{d(x,y)}(t)\right|\,dy=\int_{\mathbb{H}}\left|\int_{\mathbb{R}}\phi_{\varepsilon}(t-v)\big(\chi_{d(x,y)}(v)-\chi_{d(x,y)}(t)\big)\,dv\right|dy.$
By the triangle inequality and Fubini’s theorem, this is bounded by
$\int_{\mathbb{R}}\phi_{\varepsilon}(t-v)\int_{\mathbb{H}}\left|\chi_{d(x,y)}(v)-\chi_{d(x,y)}(t)\right|\,dy\,dv=\phi_{\varepsilon}\ast_{\mathbb{R}}A(t),$
@@
$A(v)=\int_{\mathbb{H}}|\chi_{d(x,y)}(v)-\chi_{d(x,y)}(t)|\,dy.$
By using the polar coordinates around $x$ together with the triangle inequality we have that
$A(v)\leq\int_{\mathbb{H}}\frac{\mathbf{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}}\,dy+\int_{\mathbb{H}}\frac{\mathbf{1}_{v>d(x,y)}}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\,dy<\infty.$
Hence we observe that $v\mapsto|\chi_{d(x,y)}(v)-\chi_{d(x,y)}(t)|$ has a dominating function with which the dominated convergence theorem can be used to observe that $A(v)\to 0$ as $v\to t$. $A$ is continuous, and as it is continuous, we have that
$\lim_{\varepsilon\to 0}\phi_{\varepsilon}\ast_{\mathbb{R}}A(t)=0,$
@@ 6. Geometric data
which is the geometric data appearing in the proof of Theorem1.5 in Section 8. $P_{t}$ and $P_{t,\tau}$ are the integral operators we extensively studied in the previous section, Section 5.
@@
Let $P_{t}$ and $P_{t,\tau}$ be integral operators on a compact connected hyperbolic surface $X$ with Laplace-Beltrami operator $\Delta_{X}$ such that $P_{t}$ has integral kernel
@@
and $P_{t,\tau}$ has integral kernel $K_{t,\tau}^{\Gamma}=\cos(t\tau)\,K_{t}^{\Gamma}$. Then
$\left\|\frac{1}{T}\int_{0}^{T}P_{t,\tau}aP_{t}\,dt\right\|_{HS}^{2}\leq\frac{C_{7}}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|D_{1}\setminus D(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right)$
where $\beta=\beta(\lambda_{1}(X))$, where $\lambda_{1}(X)$ is the spectral gap of $\Delta_{X}$, and where
@@
The constant appearing is
$C_{7}=\frac{7200}{\pi}.$
This is the most technical part of the whole paper. The core idea is first to write $K_{t}^{\Gamma}$ and $K_{t,\tau}^{\Gamma}$ using $K_{t}$ while simultaneously simplifying the expressions by unifying the sums over the Fuchsian group $\Gamma$ with the integrals over the fundamental domains of the surface $X$ to get integrals over the whole hyperbolic plane. This careful splitting allows us to separate a certain weight function $F_{t,t^{\prime},\rho}$ into a separate integral in which it is easier to bound. Section 7 is dedicated for the analysis of $F_{t,t^{\prime},\rho}$. Finally, we split the expression into two sides, one with the points with large injectivity radius and the other with the points with small injectivity radius. Both cases require careful analysis to acquire reasonable bounds, especially the side with the large injectivity radius for which we need the exponential mixing theorem, Theorem 1.2.
The proof relies on multiple auxiliary lemmas presented after the proof. This approach makes the proof clearer.
@@
Using that $K^{\Gamma}_{t,\tau}=\cos(t\tau)K_{t}^{\Gamma}$ and by expanding the squares, we can write the previous expression as follows:
@@
Using Fubini’s theorem to reorder the integrals yields the following equals expression:
@@
This can be written in a more compact form when gathering the factors containing $x$ into one integral and the factors containing $y$ into another integral and noticing that these two integrals are the same thus resulting into an integral squared:
@@
$\frac{1}{8\pi^{2}T^{2}}\int_{0}^{T}\int_{0}^{T}\left|\int_{D}\int_{D}a(z)a(z^{\prime})^{*}\left(\int_{\mathbb{H}}\sum_{\gamma\in\Gamma}K_{t}(x,z)K_{t^{\prime}}(x,\gamma z^{\prime})dx\right)^{2}dz^{\prime}dz\right|dt^{\prime}dt$
Coefficient $(8\,\pi^{2})^{-1}$ came from the definition of $K_{t}^{\Gamma}$:
@@
Lemma 6.3 allows us to rewrite the integral inside the absolute value of the previous integral expression in a very convenient form, eliminating but one sum. Thus the previous expression equals:
$\frac{1}{8\pi^{2}T^{2}}\int_{0}^{T}\int_{0}^{T}\left|\int_{D}\int_{\mathbb{H}}a(z)a^{*}(z^{\prime})\left(\int_{\mathbb{H}}K_{t}(x,z)K_{t^{\prime}}(x,z^{\prime})dx\right)\sum_{\gamma\in\Gamma}\left(\int_{\mathbb{H}}K_{t}(y,\gamma z)K_{t^{\prime}}(y,z^{\prime})dy\right)dz^{\prime}dz\right|dt^{\prime}dt$
which further equals
$\frac{1}{8\pi^{2}T^{2}}\int_{0}^{T}\int_{0}^{T}\left|\int_{D}a(z)\int_{\mathbb{H}}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|dt^{\prime}dt,$ (6.1)
when $F_{t,t^{\prime},\rho}$ is defined to equal
@@
where $w,w^{\prime}\in\mathbb{H}$ are any points seprated by distance $\rho$. In Section 7, where we study $F_{t,t^{\prime},\rho}$ carefully, in Lemma 7.1 we will see that the chosen points $w,w^{\prime}$ do not matter, only their distance.
@@
The definition of $F_{t,t^{\prime},d(z,z^{\prime})}$ shows that $F_{t,t^{\prime},d(z,z^{\prime})}$ equals 0 if $d(z,z^{\prime})$ is larger than $2T$, which follows from the definition of $K_{t}$ and $K_{t^{\prime}}$. Using this observation gives us the following equal expression for Expression (6.1):
$\frac{1}{8\pi^{2}T^{2}}\int_{0}^{T}\int_{0}^{T}\left|\int_{D}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|dt^{\prime}dt.$ (6.2)
Focus on the absolute values in the previous expression. They are bounded from above by
$\left|\int_{D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|$
$+\ \left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|,$ (6.3)
where the integration over $D$ is split into two: the points having injectivity radius bigger than $4T$, $D(4T)$, and the points having injecitivity radius smaller than $4T$, $D\setminus D(4T)$. The large injectivity radius case and the small injectivity radius case are bounded separately.
@@
The small injectivity radius term is bounded in Lemma 6.4. Using its bound gives the following upper bound for the previous expression:
$\|a\|_{2}^{2}\,\int_{0}^{2T}11\,e^{\beta}\,\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,(1+\rho)\,e^{-\beta\rho}d\rho$
$+\ \left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right|$
$+\ \left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|.$
In the previous, the second term is bounded by the third term as the second term has no sum and the third one does. Hence the previous expression is bounded by
$\|a\|_{2}^{2}\,\int_{0}^{2T}11\,e^{\beta}\,\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,(1+\rho)\,e^{-\beta\rho}d\rho$
+ $+\ 2\,\left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|,$
@@
where the second term is the small injectivity radius case, which is bounded in Lemma 6.5. Using this lemma gives the following upper bound for the previous expression:
@@
$\|a\|_{2}^{2}\,\int_{0}^{2T}11\,e^{\beta}\,\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,(1+\rho)\,e^{-\beta\rho}d\rho\ +\ 8\pi^{2}\,\|a\|_{\infty}^{2}|D\setminus D(4T)|\frac{e^{4T}}{\operatorname{InjRad}_{X}}\int_{0}^{2T}\sinh(\rho)F_{t,t^{\prime},\rho}^{2}d\rho.$ (6.4)
The latter term is bounded by
$8\pi^{2}\,\|a\|_{\infty}^{2}|D\setminus D(4T)|\frac{e^{4T}}{\operatorname{InjRad}_{X}}\int_{0}^{2T}11e^{\beta}\sinh(\rho)(1+\rho)e^{-\beta\rho}e^{2\beta T}F_{t,t^{\prime},\rho}^{2}d\rho,$
since
$11\,e^{\beta}(1+\rho)e^{-\beta\rho}e^{\beta\,2T}\geq 1.$
Hence Expression (6.4) is bounded by:
@@
$11e^{\beta}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\frac{e^{4T+2T\beta}}{\operatorname{InjRad}_{X}}\right)\int_{0}^{2T}\sinh(\rho)(1+\rho)F_{t,t^{\prime},\rho}^{2}d\rho.$
We have now acquired good bounds for the absolute values appearing in Expression (6.1), and this expression is thus bounded by
@@
$\frac{11}{8\pi^{2}T^{2}}\,e^{\beta}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\frac{e^{4T+2T\beta}}{\operatorname{InjRad}_{X}}\right)\int_{0}^{T}\int_{0}^{T}\int_{0}^{2T}\sinh(\rho)(1+\rho)F_{t,t^{\prime},\rho}^{2}d\rho dt^{\prime}dt.$
As $F_{t,t^{\prime},\rho}$ equals $F_{t^{\prime},t,\rho}$ and since we integrate from 0 to $T$ both $t$ and $t^{\prime}$, we can use symmetry to integrate the inner integral of $t^{\prime}$ from 0 to only $t$ twice. Hence we write the previous expression as
@@
$\frac{22}{8\pi^{2}T^{2}}\,e^{\beta}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\frac{e^{4T+2T\beta}}{\operatorname{InjRad}_{X}}\right)\int_{0}^{T}\int_{0}^{t}\int_{0}^{2T}\sinh(\rho)(1+\rho)F_{t,t^{\prime},\rho}^{2}d\rho dt^{\prime}dt.$ (6.5)
The innermost integral can be bounded using Proposition 6.6, which uses the bound acquired for $F_{t,t^{\prime},\rho}$ in Section 7. We get the following upper bound for the previous expression:
@@
$\frac{22}{8\pi^{2}T^{2}}\,e^{\beta}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\frac{e^{4T+2T\beta}}{\operatorname{InjRad}_{X}}\right)\int_{0}^{T}\int_{0}^{t}\frac{C_{6}}{\beta^{2}}\,e^{-\frac{\beta}{2}(t-t^{\prime})}dt^{\prime}dt,$
where
$C_{6}=480\pi.$
The remaining two integrals have
$\frac{2T\,C_{6}}{\beta^{3}}$
as an upper bound, which is seen after evaluating these integrals. Then the previous integral expression has the following upper bound:
@@
$\frac{C_{7}}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
where
$\frac{22\cdot 2\cdot C_{6}\cdot e}{8\pi^{2}}\leq\frac{7200}{\pi}=C_{7},$
where we used that $e^{\beta}$ is less than $e$. This is also an upper bound for the Hilbert-Schmidt norm we set out to bound. It is of the wanted form which concludes the proof.
@@
The remaining of this section contains the lemmas needed to prove Proposition 6.1 together with the contents of Section 7. The following lemma allows us to write the integration containing the kernels living on the hyperbolic surface using the kernels living on the hyperbolic plane.
@@
$\int_{D}K^{\Gamma}(x,z)L^{\Gamma}(x,z^{\prime})dx=\int_{\mathbb{H}}\sum_{\gamma\in\Gamma}K(x,z)L(x,\gamma_{2}z^{\prime})dx,$
@@
The proof is on uniting the sum over the Fuchsian group and the integration over the fundamental domain into an integration over the hyperbolic plane.
@@
by opening the kernels using their sum form. Doing a change of variables $\gamma_{2}\mapsto\gamma_{1}\gamma_{2}$ for the sum allows us to write the previous expression as
@@
Uniting the integral and the sum of $\gamma_{1}$ into the integral over the universal cover gives us the following equal expression:
@@
where we opted to use $\gamma$ since there is no need to distinguish the sums anymore. This concludes the proof.
The following lemma allows us to simplify the notation by removing one sum over the Fuchsian group.
@@
Let $K$ and $L$ be integral kernels on compact connected hyperbolic surface $X=\Gamma\setminus\mathbb{H}$. Let $a$ be a function on $X$. Then:
$\int_{D}a^{*}(z^{\prime})\left(\int_{\mathbb{H}}\sum_{\gamma\in\Gamma}K(x,z)L(x,\gamma z^{\prime})dx\right)^{2}dz^{\prime}$
$=\qquad$ $\int_{\mathbb{H}}a^{*}(z^{\prime})\left(\int_{\mathbb{H}}K(x,z)L(x,z^{\prime})dx\right)\sum_{\gamma\in\Gamma}\left(\int_{\mathbb{H}}K(y,\gamma z)L(y,z^{\prime})dy\right)dz^{\prime},$
@@
The proof is based on the same principle as in the proof of Lemma 6.2: we unite the sum over the Fuchsian group and the integral over the fundamental domain into an integration over the hyperbolic plane.
@@
$\int_{D}a^{*}(z^{\prime})\left(\int_{\mathbb{H}}\sum_{\gamma\in\Gamma}K(x,z)L(x,\gamma z^{\prime})dx\right)^{2}dz^{\prime}.$
@@
$\int_{D}\sum_{\gamma_{1}\in\Gamma}\sum_{\gamma_{2}\in\Gamma}a^{*}(z^{\prime})\left(\int_{\mathbb{H}}K(x,z)L(x,\gamma_{1}z^{\prime})dx\right)\left(\int_{\mathbb{H}}K(y,z)L(y,\gamma_{2}z^{\prime})dy\right)dz^{\prime}.$
Doing a change of variables $\gamma_{2}\mapsto\gamma_{2}^{-1}\gamma_{1}$ in the sum allows us to write the previous expression as
$\int_{D}\sum_{\gamma_{1}\in\Gamma}\sum_{\gamma_{2}\in\Gamma}a^{*}(z^{\prime})\left(\int_{\mathbb{H}}K(x,z)L(x,\gamma_{1}z^{\prime})dx\right)\left(\int_{\mathbb{H}}K(y,z)L(y,\gamma_{2}^{-1}\gamma_{1}z^{\prime})dy\right)dz^{\prime},$
@@
$\int_{D}\sum_{\gamma_{1}\in\Gamma}\sum_{\gamma_{2}\in\Gamma}a^{*}(z^{\prime})\left(\int_{\mathbb{H}}K(x,z)L(x,\gamma_{1}z^{\prime})dx\right)\left(\int_{\mathbb{H}}K(y,z)L(\gamma_{2}y,\gamma_{1}z^{\prime})dy\right)dz^{\prime},$
when using that $L(x,\gamma y)=L(\gamma^{-1}x,y)$. Uniting the integral and the sum of $\gamma_{1}$ into the integral over the universal cover is possible since $a^{*}$ can be seen as $\Gamma$-periodic function on $\mathbb{H}$. This gives the following equal expression:
$\int_{\mathbb{H}}\sum_{\gamma\in\Gamma}a^{*}(z^{\prime})\left(\int_{\mathbb{H}}K(x,z)L(x,z^{\prime})dx\right)\left(\int_{\mathbb{H}}K(y,z)L(\gamma y,z^{\prime})dy\right)dz^{\prime},$ (6.6)

### v2

@@ 5. Propagators
We want to show that the following expression goes to 0 as $\varepsilon\to 0$:
$\int_{\mathbb{H}}\left|\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)-\chi_{d(x,y)}(t)\right|\,dy=\int_{\mathbb{H}}\left|\int_{\mathbb{R}}\phi_{\varepsilon}(t-v)\big(\chi_{d(x,y)}(v)-\chi_{d(x,y)}(t)\big)\,dv\right|dy.$ (5.2)
Here, we used that the integral over the Gaussian mollifier equals $1$:
$\chi_{d(x,y)}(t)=\int_{\mathbb{R}}\phi_{\varepsilon}(t-v)\,\chi_{d(x,y)}(t)\,dv.$
By the triangle inequality and Fubini’s theorem, Expression (5.2) is bounded by
$\int_{\mathbb{R}}\phi_{\varepsilon}(t-v)\int_{\mathbb{H}}\left|\chi_{d(x,y)}(v)-\chi_{d(x,y)}(t)\right|\,dy\,dv=\phi_{\varepsilon}\ast_{\mathbb{R}}A_{t}(t),$
@@
$A_{t}(v)=\int_{\mathbb{H}}|\chi_{d(x,y)}(v)-\chi_{d(x,y)}(t)|\,dy.$
We see that $A_{t}(t)=0$, and we want to show that $\lim_{v\to t}A_{t}(v)=0$. By the triangle inequality:
$A_{t}(v)\leq\int_{\mathbb{H}}\frac{\mathbf{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}}\,dy+\int_{\mathbb{H}}\frac{\mathbf{1}_{v>d(x,y)}}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\,dy<\infty.$
The finiteness can be verified using polar coordinates around $x$. Hence, $y\mapsto|\chi_{d(x,y)}(v)-\chi_{d(x,y)}(t)|$ has a dominating function when $v$ is in a compact interval around $t$. By the dominated convergence theorem, $A_{t}(v)\to 0$ as $v\to t$. Since $A_{t}(t)=0$ and $A_{t}$ is continuous at $t$, the properties of convolution yield
$\lim_{\varepsilon\to 0}\phi_{\varepsilon}\ast_{\mathbb{R}}A_{t}(t)=0,$
@@ 6. Geometric data
which is the geometric data appearing in the proof of Theorem 1.5 in Section 8. $P_{t}$ and $P_{t,\tau}$ are the integral operators we extensively studied in the previous section, Section 5. We now state the proposition.
@@
Let $P_{t}$ and $P_{t,\tau}$ be integral operators on a compact connected hyperbolic surface $X=\Gamma\setminus\mathbb{H}$ with Laplace–Beltrami operator $\Delta_{X}$ such that $P_{t}$ has integral kernel $K_{t}^{\Gamma}$, defined as
@@
and $P_{t,\tau}$ has integral kernel $K_{t,\tau}^{\Gamma}=\cos(t\tau)\,K_{t}^{\Gamma}$. Let $a$ be a function on $X$ with mean zero. Then
$\left\|\frac{1}{T}\int_{0}^{T}P_{t,\tau}^{*}aP_{t}\,dt\right\|_{HS}^{2}\leq\frac{C_{7}}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,\|a\|_{\infty}^{2}\,|X\setminus X(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right)$
where $\beta=\beta(\lambda_{1}(X))$, $|X\setminus X(4T)|$ is the volume of the points of $X$ having injectivity radius less than $4T$, $\lambda_{1}(X)$ is the spectral gap of $\Delta_{X}$, and where
@@
The constant appearing is $C_{7}=80000$.
By careful bounding and rewriting the Hilbert-Schmidt norm expression, we get a tidy expression in which the observable $a$ is separated from the geometric data arising from the choice of the propagator; this geometric data is contained in a certain weight function $F_{t,t^{\prime},\rho}$. This weight is studied in Section 7. The bound is acquired by separating an integration over the fundamental domain into a part dealing with the points with large injectivity radius and a part dealing with the points with small injectivity radius. In the large injectivity radius case, the exponential mixing theorem, Theorem 1.2, gives a crucial decay needed to acquire the bounds. The final bounds are obtained by using bounds for $F_{t,t^{\prime},\rho}$.
@@
We also used that $P_{t,\tau}$ is self-adjoint. Using that $K^{\Gamma}_{t,\tau}=\cos(t\tau)K_{t}^{\Gamma}$ and by expanding the squares, we can write the previous expression as follows:
@@
Using Fubini’s theorem to reorder the integrals yields the following equal expression:
@@
The cosines were bounded from above by $1$ in the previous step. The previous expression can be written in a more compact form when gathering the factors containing $x$ into one integral and the factors containing $y$ into another integral and noticing that these two integrals are the same, thus resulting in an integral squared:
@@
$\frac{1}{8\pi^{2}T^{2}}\int_{0}^{T}\int_{0}^{T}\left|\int_{D}\int_{D}a(z)a(z^{\prime})^{*}\left(\int_{\mathbb{H}}\sum_{\gamma\in\Gamma}K_{t}(x,z)K_{t^{\prime}}(x,\gamma z^{\prime})dx\right)^{2}dz^{\prime}dz\right|dt^{\prime}dt,$
where the coefficient $(8\,\pi^{2})^{-1}$ comes from the definition of $K_{t}^{\Gamma}$:
@@
Lemma 6.3 allows us to rewrite the integral inside the absolute value in a convenient form, eliminating all but one sum. Thus the previous expression equals:
$\frac{1}{8\pi^{2}T^{2}}\int_{0}^{T}\int_{0}^{T}\left|\int_{D}a(z)\int_{\mathbb{H}}a(z^{\prime})^{*}F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|dt^{\prime}dt,$ (6.1)
where $F_{t,t^{\prime},\rho}$ is defined as
@@
with $w,w^{\prime}\in\mathbb{H}$ separated by distance $\rho$; Lemma 7.1 ensures that $F_{t,t^{\prime},\rho}$ is well-defined.
@@
The definition of $F_{t,t^{\prime},d(z,z^{\prime})}$ shows that $F_{t,t^{\prime},d(z,z^{\prime})}=0$ if $d(z,z^{\prime})>2T$, which follows from the definitions of $K_{t}$ and $K_{t^{\prime}}$. Using this, Expression (6.1) can be rewritten by restricting the integration domain of $z^{\prime}$ as follows:
$\frac{1}{8\pi^{2}T^{2}}\int_{0}^{T}\int_{0}^{T}\left|\int_{D}a(z)\int_{B(z,2T)}a(z^{\prime})^{*}F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|dt^{\prime}dt.$ (6.2)
We split the integral over $D$ into points with injectivity radius greater than $4T$, $D(4T)$, and points with injectivity radius less than $4T$, $D\setminus D(4T)$:
$\left|\int_{D}\cdots\right|\leq\left|\int_{D(4T)}\cdots\right|+\left|\int_{D\setminus D(4T)}\cdots\right|.$
By moving the absolute value inside the latter term and bounding $|a|$ by its supremum norm, the previous expression yields the following upper bound:
$\left|\int_{D(4T)}a(z)\int_{B(z,2T)}a(z^{\prime})^{*}F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|$
[…]
where the large and small injectivity radius terms are bounded separately.
@@
The small injectivity radius term is bounded in Lemma 6.4. Using this estimate, we obtain the following upper bound:
$30\|a\|_{2}^{2}\,\int_{0}^{2T}\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,(1+\rho)\,e^{-\beta\rho}d\rho+2\,\|a\|_{\infty}^{2}\int_{D\setminus D(4T)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz.$
@@
The second term of the previous expression is bounded in Lemma 6.5, yielding the following upper bound:
@@
$30\|a\|_{2}^{2}\,\int_{0}^{2T}\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,(1+\rho)\,e^{-\beta\rho}d\rho\ +\ 8\pi^{2}\,\|a\|_{\infty}^{2}|D\setminus D(4T)|\frac{e^{4T}}{\operatorname{InjRad}_{X}}\int_{0}^{2T}\sinh(\rho)F_{t,t^{\prime},\rho}^{2}d\rho.$
This is further bounded by
@@
$8\pi^{2}\left(\|a\|_{2}^{2}\,+\,\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\,\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right)\int_{0}^{2T}\sinh(\rho)e^{-\beta\rho}(1+\rho)F_{t,t^{\prime},\rho}^{2}d\rho,$
using that $8\pi^{2}>30$ and $e^{2T}\,e^{-\beta\rho}\,(1+\rho)\geq 1$.
Hence, Expression (6.1) is bounded by
@@
$\frac{1}{T^{2}}\,\left(\|a\|_{2}^{2}\,+\,\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right)\int_{0}^{T}\int_{0}^{T}\int_{0}^{2T}\sinh(\rho)e^{-\beta\rho}(1+\rho)F_{t,t^{\prime},\rho}^{2}d\rho dt^{\prime}dt.$
By symmetry of $F_{t,t^{\prime},\rho}$ in $t$ and $t^{\prime}$, we can restrict the inner integral to $0\leq t^{\prime}\leq t$ and multiply by $2$:
@@
$\frac{2}{T^{2}}\,\left(\|a\|_{2}^{2}\,+\,\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right)\int_{0}^{T}\int_{0}^{t}\int_{0}^{2T}\sinh(\rho)e^{-\beta\rho}(1+\rho)F_{t,t^{\prime},\rho}^{2}d\rho dt^{\prime}dt.$
Applying Proposition 6.6 to the innermost integral, we obtain the upper bound
@@
$\frac{2}{T^{2}}\left(\|a\|_{2}^{2}+\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right)\int_{0}^{T}\int_{0}^{t}\frac{C_{6}}{\beta^{2}}\,e^{-\frac{\beta}{4}(t-t^{\prime})}\,dt^{\prime}\,dt,$
where $C_{6}=10000$. Integrating in $t^{\prime}$ and $t$ gives an upper bound of $4TC_{6}/\beta^{3}$, so the entire expression is bounded by
@@
$\frac{C_{7}}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
with $2\cdot 4\cdot C_{6}=80000=C_{7}$.
@@
The remaining part of this section contains the lemmas needed to prove Proposition 6.1, together with the contents of Section 7. The following lemma allows us to write the integration containing the kernels living on the hyperbolic surface using the kernels living on the hyperbolic plane.
@@
$\int_{D}K^{\Gamma}(x,z)L^{\Gamma}(x,z^{\prime})dx=\int_{\mathbb{H}}\sum_{\gamma\in\Gamma}K(x,z)L(x,\gamma z^{\prime})dx,$
@@
The proof is based on uniting the sum over the group $\Gamma$ and the integration over the fundamental domain $D$ into an integration over the hyperbolic plane.
@@
by opening the kernels using their sum forms. Doing a change of variables $\gamma_{2}\mapsto\gamma_{1}\gamma_{2}$ for the inner sum allows us to write the previous expression as
@@
Uniting the integral and the sum over $\gamma_{1}$ into the integral over the universal cover gives us the following equal expression:
@@
where we opted to use $\gamma$ since there is no need to distinguish the sums anymore. This concludes the proof. ∎
The following lemma allows us to simplify an integral expression in which sums over the group $\Gamma$ appear.
@@
Let
be integral kernels on a compact connected hyperbolic surface $X=\Gamma\setminus\mathbb{H}$, where $K$ and $L$ are integral kernels on $\mathbb{H}$ that depend only on distance $d(x,y)$. Let $a$ be a function on $X$. Then
$\int_{D}a(z^{\prime})^{*}\left(\int_{\mathbb{H}}\sum_{\gamma\in\Gamma}K(x,z)L(x,\gamma z^{\prime})dx\right)^{2}dz^{\prime}$
$=\ \int_{\mathbb{H}}a(z^{\prime})^{*}\left(\int_{\mathbb{H}}K(x,z)L(x,z^{\prime})dx\right)\sum_{\gamma\in\Gamma}\left(\int_{\mathbb{H}}K(y,\gamma z)L(y,z^{\prime})dy\right)dz^{\prime},$
@@
The proof is based on the same principle as the proof of Lemma 6.2: we unite the sum over the group and the integral over the fundamental domain into an integration over the hyperbolic plane.
@@
$\int_{D}a(z^{\prime})^{*}\left(\int_{\mathbb{H}}\sum_{\gamma\in\Gamma}K(x,z)L(x,\gamma z^{\prime})dx\right)^{2}dz^{\prime}.$
@@
$\int_{D}\sum_{\gamma_{1}\in\Gamma}\sum_{\gamma_{2}\in\Gamma}a(z^{\prime})^{*}\left(\int_{\mathbb{H}}K(x,z)L(x,\gamma_{1}z^{\prime})dx\right)\left(\int_{\mathbb{H}}K(y,z)L(y,\gamma_{2}z^{\prime})dy\right)dz^{\prime}.$
Doing a change of variables $\gamma_{2}\mapsto\gamma_{2}^{-1}\gamma_{1}$ in the inner sum allows us to write the previous expression as
$\int_{D}\sum_{\gamma_{1}\in\Gamma}\sum_{\gamma_{2}\in\Gamma}a(z^{\prime})^{*}\left(\int_{\mathbb{H}}K(x,z)L(x,\gamma_{1}z^{\prime})dx\right)\left(\int_{\mathbb{H}}K(y,z)L(y,\gamma_{2}^{-1}\gamma_{1}z^{\prime})dy\right)dz^{\prime},$
@@
$\int_{D}\sum_{\gamma_{1}\in\Gamma}\sum_{\gamma_{2}\in\Gamma}a(z^{\prime})^{*}\left(\int_{\mathbb{H}}K(x,z)L(x,\gamma_{1}z^{\prime})dx\right)\left(\int_{\mathbb{H}}K(y,z)L(\gamma_{2}y,\gamma_{1}z^{\prime})dy\right)dz^{\prime},$
when using that $L(x,\gamma y)=L(\gamma^{-1}x,y)$. Uniting the integral and the sum over $\gamma_{1}$ into the integral over the universal cover is possible since $a^{*}$ can be seen as a $\Gamma$-periodic function on $\mathbb{H}$. This gives the following equal expression:
$\int_{\mathbb{H}}\sum_{\gamma\in\Gamma}a(z^{\prime})^{*}\left(\int_{\mathbb{H}}K(x,z)L(x,z^{\prime})dx\right)\left(\int_{\mathbb{H}}K(y,z)L(\gamma y,z^{\prime})dy\right)dz^{\prime},$ (6.4)